Your piggy bank contains $19.75 in dimes and quarters. There are 100 coins in all. How many dimes are there?
Answer by mananth(16946)

let the quarters be x
the dimes will be 100-x
..
the value of both = 1975 cents
25x+10*(100-x)=1975
25x+1000-10x=1975
15x+1000=1975
add-1000
15x+1000-1000=1975-1000
15x=975
x=975/15
number of quarters = 65
so dimes will be 35